OpenStudy (anonymous):

Solve the equation: 9x − 2 = 4(x + 3)

OpenStudy (anonymous):

5x=14 x=14/5

Parth (parthkohli):

Distribute on the right hand side. \(\Large \color{Black}{\Rightarrow 9x - 2 = 4x + 12 }\) Add 2 to both sides. \(\Large \color{Black}{\Rightarrow 9x = 4x + 14 }\) Subtract 4x from both sides. \(\Large \color{Black}{\Rightarrow 5x = 14}\) Divide both sides by 5. \(\Large \color{Black}{\Rightarrow x = 14/5 = 2.8 }\)